# Free Record-Level Privacy Risk Evaluation Through Artifact-Based Methods

### v1

@@ Abstract
Membership inference attacks (MIAs) are widely used to empirically assess the privacy risks of samples used to train a target machine learning model. State-of-the-art methods however require training hundreds of shadow models, with the same size and architecture of the target model, solely to evaluate the privacy risk. While one might be able to afford this for small models, the cost often becomes prohibitive for medium and large models.
We here instead propose a novel approach to identify the at-risk samples using only artifacts available during training, with little to no additional computational overhead. Our method analyzes individual per-sample loss traces and uses them to identify the vulnerable data samples. We demonstrate the effectiveness of our artifact-based approach through experiments on the CIFAR10 dataset, showing high precision in identifying vulnerable samples as determined by a SOTA shadow model-based MIA (LiRA [7]). Impressively, our method reaches the same precision as another SOTA MIA when measured against LiRA, despite it being orders of magnitude cheaper. We then show LT-IQR to outperform alternative loss aggregation methods, perform ablation studies on hyperparameters, and validate the robustness of our method to the target metric. Finally, we study the evolution of the vulnerability score distribution throughout training as a metric for model-level risk assessment.
@@ 1 Introduction
Figure 1: Precision@k=1% (250 samples) when identifying vulnerable samples determined by LiRA attack at a variable FPR threshold (log scale).
Machine Learning (ML) models are increasingly trained or fine-tuned by companies, governments, and academic researchers often using private, sensitive, or confidential data. This includes medical [17], legal [11], and financial data [3].
@@
Membership inference attacks (MIAs) are widely used to empirically evaluate whether an ML model has memorized individual records and the overall privacy risk releasing the model poses. ML models have indeed been shown to memorize their training data [35, 8, 7], allowing an adversary to detect the presence of a certain data sample in the training dataset [41, 34, 7], or even reconstruct the entire training samples with little prior knowledge [6, 14].
Model developers often rely on running empirical privacy attacks against their own models to assess potential risks associated with the public release of these models or their predictions [31]. This empirical approach to privacy analysis is often used alongside the privacy-preserving training methods with formal guarantees, such as DP-SGD [1], or as a standalone mitigation where the severe impact of differentially private training on model utility makes it impractical [31].
@@
State-of-the-art privacy attacks are, however, computationally expensive to run. They require training multiple, often hundreds, of shadow models [7, 40, 45], with each shadow model requiring the same amount of computational resources as one original training run. Depending on the scale of the target model and the training dataset, this renders practical uses of such attacks challenging–e.g. as part of the selection process of architectures, training parameters, etc–if not completely infeasible - for example, in the case of Large Language Models.
Contribution. We here propose a new class of methods to estimate a model’s vulnerability to privacy attacks exclusively utilizing artifacts available for free at training time. Artifact-based methods allow model developers to leverage their white-box access to the model throughout the training to identify data samples vulnerable to privacy attacks with little to no computational overhead.
To first underscore the difficulty of the task, we study the level of agreement between SOTA MIA methods and show them to only partially agree on the most vulnerable points. Despite assuming the same attacker knowledge and being similar shadow model-based methods, LiRA and Attack R only partially agree on the most vulnerable points.
We then consider a binary classification setup, defining vulnerable samples as the ones successfully identified as members by the Likelihood Ratio Attack (LiRA) attack [7] at a fixed low False Positive Rate (FPR), a highly imbalanced task.
We then propose Loss Trace IQR (LT-IQR), as a simple yet robust aggregation score for sample-level loss. In a series of experiments on the CIFAR10 [19] dataset, we confirm our proposed method to reach a level of agreement on par with computationally expensive SOTA shadow model-based MIAs despite being essentially free. LT-IQR (75%-25%), for instance, achieves 61% Precision@k=1% in predicting vulnerable samples defined by LiRA at a low FPR threshold of $10^{-3}$, a result in line with Attack R [40]. This demonstrates that our method is as good at identifying vulnerable samples as a state-of-the-art shadow-model-based MIA, while being essentially free for the model developer.
Additionally, we compare LT-IQR to multiple potential aggregation methods for the loss trace and show it to outperform these. We perform ablation studies on the choice of quantile parameters in LT-IQR, and validate the robustness of our method to the target metric. Finally, we study the evolution of the vulnerability score distribution across the training process as a metric for model-level risk assessment.
@@
Our cheap yet effective approach enables model developers to assess the privacy risks associated with a given model release, and deploy appropriate mitigations if needed, including unlearning, data removal, or retraining with differential privacy. Given its low computational cost, our method could be especially useful to compare the privacy leakage of different hyperparameter setups, or as an early indication to run a full privacy attack on a subset of candidate models.
@@ 2.1 Membership Inference Attacks
Membership Inference Attacks (MIAs) aim to determine whether a certain data sample was used to train a given model. They typically rely on a strong yet realistic attacker [31, 28] and are used as an upper bound for other attacks including extraction  [10, 30, 25] and attribute inference attacks.
First introduced in the context of ML training by Shokri et al. [34], current MIAs against ML models typically assume a black-box setting and rely on the loss of the trained model as the main signal for memorization [34, 41, 7, 40, 45]. Sablayrolles et al. [33], for instance, argue that an optimal attack only needs to depend on the loss and that white-box access would give no additional advantage to an attacker.
@@
While attacks have been developed with lower computational costs [41, 45], they either suffer from low utility compared to the shadow-model-based attacks [41], or still require training some shadow models [45], an expensive task, with the attack performance being positively correlated with the number of shadow models.
@@ 2.2 Individual vulnerability
Thanks to the extensive literature on the topic, vulnerability to privacy attacks is now understood to be highly non-uniform across data distributions and is measured via TPR@low FPR. Empirical privacy risk is indeed typically concentrated around a small number of data outliers [13, 5, 9]. Additionally, recent works [2, 15] argued that common MIA evaluations fail to capture the risk for the most vulnerable samples. Throughout this work we focus on record-specific risk, with the main goal to identify the samples most vulnerable to privacy attacks.
While most MIAs are postulated as a privacy game with a fixed model [34, 41, 7], Ye et al. [40] have adapted the definition to use a fixed record instead, focusing on individual vulnerability. Subsequent work has developed record-specific attacks [4] and evaluated existing techniques under the new definition [15, 20]. In particular, Leemann et al. [20] investigate the connection between MIA vulnerability and certain per-instance properties: final loss, parameter gradients, local loss Hessian, and others, identifying the final loss to be the best predictor of the MIA performance. Additionally, Tobaben et al. [38] study the impact of the average gradient norm throughout training on the MIA performance and find a strong correlation across multiple datasets. Unlike our work, however, both Leemann et al. [20] and Tobaben et al. [38] study the correlation between the data properties and attack success while we focus on using training artifacts to accurately identify the few most vulnerable samples.
@@ 2.3 Tracing memorization
Unlike privacy analyses described above, we consider an artifact-based approach where an attacker, here the model developer, has access to all intermediate model states during training. This setup, for instance, allows us to use the loss throughout training as an additional source of information on top of the commonly used final loss.
@@
Some prior works have considered tracing privacy metrics throughout training. Pruthi et al. [32] tracks the loss before and after each training step to estimate the influence of a training sample on model’s predictions. Lesci et al. [21] characterize a model’s counterfactual memorization [46] by observing losses on a subset of data throughout training. Liu et al. [23] and Li et al. [22] propose using loss trajectory in a traditional MIA setup by tracking the loss during the target model distillation process.
@@
For differentially private training, privacy auditing often adopts a similar threat model to us. Privacy auditing has emerged as a way to either assess the tightness of the theoretical privacy analysis [16, 30, 44] or verify the correctness of DP algorithm implementations [39, 29]. Privacy auditing empirically estimates the lower bound for the differential privacy guarantees of the model training process by running privacy attacks and measuring their success. Assumed threat models can range from a weak adversary with only black-box access to the model [16, 30, 36], to a white-box adversary capable of observing intermediate model updates [30, 29, 36], to a very strong adversary with full control over the training dataset apart from the target record [30]. Prior works on privacy auditing however typically rely on data poisoning attacks, instantiating the MIAs against data samples (canaries) purposely injected into the training dataset.
@@
Proposed attack methods vary depending on the threat model. For the black-box attacks, a simple loss threshold [41] is typically used. For white-box adversaries with access to the intermediate model states, a threat model similar to the one we assume in this work, prior works suggested using aggregate metrics over the entire training process. Proposed methods include taking the maximum or average of the loss [30], or computing the dot product between the canary gradient and the overall batch update [29, 36].
@@ 3.1.1 Model training
Let $f_{\theta}:\mathcal{X}\rightarrow[0,1]^{N}$ be a machine learning classification model with an input domain $\mathcal{X}$ and an output domain $\mathcal{Y}$ with $N=|\mathcal{Y}|$ classes, parametrized by a parameter vector $\theta$. $f_{\theta}(x)$ outputs a probability distribution over $N$ classes. For brevity, we denote the probability of a given class $y$ as $f_{\theta}(x)_{y}$. We denote train and test datasets as $D_{\text{train}}=\{(x_{i},y_{i})\mid x_{i}\in\mathcal{X},y_{i}\in\mathcal{Y}\}$ and $D_{\text{test}}=\{(x^{*}_{i},y^{*}_{i})\mid x^{*}_{i}\in\mathcal{X},y^{*}_{i}\in\mathcal{Y}\}$ respectively. Let $\mathcal{T}$ be a training process such that the final model weights $\theta$ are obtained as $\theta\leftarrow\mathcal{T}(D_{\text{train}})$. We consider the training process to be performed over $S$ epochs, and additionally denote intermediate model state after $s=1\dots S$ epochs as $\theta_{s}\leftarrow\mathcal{T}_{s}(D_{\text{train}})$. Note that $\theta_{0}$ is the initial model state and $\theta_{S}=\theta$ is the parameter vector of a fully trained model.
@@
We denote the loss on a sample $(x,y)$ after $s$ training epochs as $\ell_{s}(x,y)=\ell(f_{\theta_{s}}(x),y)$. We also define $\mathcal{L}_{\text{train}}^{s}$ and $\mathcal{L}_{\text{test}}^{s}$ as average losses at epoch $s$ on train and test sets respectively.
@@
3.1.2 Membership Inference Attacks
A membership inference attack (MIA) aims to determine whether a given data sample was used to train a model. Formally, we define a membership inference attack as a function $A:\mathcal{X}\times\mathcal{Y}\times\Theta\rightarrow[0,1]$, where $A(x,y,\theta)=1$ indicates that the attacker believes the sample $(x,y)$ was used to train the model with parameters $\theta$, and $A(x,y,\theta)=0$ indicates when they believe it was not. In particular, we consider SOTA MIAs that output a confidence score, where higher values indicate stronger belief that the sample was used during training. The binary classification decision is then made by comparing this score against a threshold $\tau$:

### v3

@@ Abstract
Membership inference attacks (MIAs) are widely used to empirically assess privacy risks in machine learning models, both providing model-level vulnerability metrics and identifying the most vulnerable training samples. State-of-the-art methods, however, require training hundreds of shadow models with the same architecture as the target model. This makes the computational cost of assessing the privacy of models prohibitive for many practical applications, particularly when used iteratively as part of the model development process and for large models. We propose a novel approach for identifying the training samples most vulnerable to membership inference attacks by analyzing artifacts naturally available during the training process. Our method, Loss Trace Interquartile Range (LT-IQR), analyzes per-sample loss trajectories collected during model training to identify high-risk samples without requiring any additional model training. Through experiments on standard benchmarks, we demonstrate that LT-IQR achieves 92% precision@k=$1\%$ in identifying the samples most vulnerable to state-of-the-art MIAs. This result holds across datasets and model architectures with LT-IQR outperforming both traditional vulnerability metrics, such as loss, and lightweight MIAs using few shadow models. We also show LT-IQR to accurately identify points vulnerable to multiple MIA methods and perform ablation studies. We believe LT-IQR enables model developers to identify vulnerable training samples, for free, as part of the model development process. Our results emphasize the potential of artifact-based methods to efficiently evaluate privacy risks.
@@ 1 Introduction
Figure 1: Precision@k=1% (250 samples) when identifying vulnerable samples determined by LiRA attack at a variable FPR threshold
In recent years, machine learning (ML) models have become increasingly prevalent within sensitive domains. Companies, governments, and academic researcher are now routinely training and fine-tuning these models using confidential or sensitive data from areas such as the healthcare [17], legal [9], and financial [2] sectors.
@@
ML models have been shown to memorize their training data [40, 6, 5, 30], potentially allowing an adversary to detect the presence of a certain data sample in the training dataset [46, 39, 5] and, in some cases, to even reconstruct training samples [4, 14].
Membership inference attacks (MIAs) are widely used by model developers to empirically evaluate the privacy risk associated with the release of the model [46, 5, 45]. Running an MIA against their own model allows developers to learn two valuable insights. First, it provides a sense of the model-level risk, the number of points an adversary might be able to confidently identify as members of the training dataset typically measured as True Positive Rate (TPR) at a certain False Positive Rate (FPR) threshold. Second, it identifies the points that are most likely to be vulnerable to an attack - e.g. the training data points where the attack confidence is above the threshold for a given FPR.
Model developers will typically evaluate the overall privacy risk by taking into consideration the estimated model-level risk (e.g. TPR@low FPR), the sensitivity of the data, the anticipated threat model (white-box or black-box) and model deployment strategy, as well as the analysis of the specific points that might be a risk (e.g. to ensure there is no disparate impact). Based on these, the developer would determine if releasing the model is deemed safe from a privacy perspective. As such, vulnerability to MIAs is an important factor in determining whether a model is considered anonymous including from a legal perspective as emphasized in the opinion published by the European Data Protection Board on AI last month [12].
When the risk of the utility-maximizing model–the model developed without taking privacy considerations into account–is deemed too high, defenses will be put in place. This includes both privacy-preserving training methods using formal guarantees, such as DP-SGD [1], as well as approaches–including approaches targeted at specific points–such as increased regularization, reducing the model size, removing the most vulnerable points from the training data (which has been showed to decrease the overall risk [7]) or unlearning them from the trained model. Knowing both which and how many points are vulnerable to an MIA is essential to the development and evaluation of defense strategies.
@@
State-of-the-art privacy attacks are, however, computationally expensive to run. They require training multiple, often hundreds, of shadow models [5, 45, 51], with each shadow model requiring the same amount of computational resources as one original training run. Their computational complexity increases with model and dataset scale, making such attacks impractical if not infeasible—particularly for large language models—and significantly constraining their utility in processes like architecture or training parameter selection.
Contribution. We here propose a new class of artifact-based methods to estimate privacy risks which we instantiate on the task of detecting points that might be vulnerable to MIAs. Our method leverages access to training artifacts to detect the training samples most at risk of being identified as members by state-of-the-art MIAs. We define a training data point as vulnerable if a given MIA confidently and correctly predicts its membership, i.e. the point is predicted as a member at a fixed low False Positive Rate (FPR).
As our main contribution, we propose Loss Trace Interquartile Range (LT-IQR), a novel method analyzing the per-sample loss dynamics throughout training to identify the most vulnerable training samples. It is based on the intuition that points prone to be memorized are likely to be outliers and therefore harder for the model to learn. This would lead to the loss of these samples converging slower than that of an average point. Specifically, LT-IQR computes the interquartile range across intermediate losses collected once per epoch, and is shown to be a good predictor of individual vulnerability.
LT-IQR achieves 92% Precision@k=1% in predicting training samples from CIFAR-10 [19] vulnerable to Likelihood Ratio Attack (LiRA) [5] at a low FPR threshold of $10^{-3}$, significantly outperforming traditional vulnerability predictors based on metrics like final loss or gradient norms. This advantage holds across datasets (CIFAR-10, CIFAR-100, and CINIC-10) and model architectures (ResNet-20, WideResNet28-2, WideResNet40-4). Notably, LT-IQR even outperforms lightweight membership inference attacks that require training shadow models (like RMIA [51] with 2 shadow models) when we use the same Precision@k metric to measure the agreement between the attacks.
While initially focused on predicting samples vulnerable specifically to the LiRA attack, we later expand the definition and focus on more generic vulnerability to any of the widely adopted state-of-the-art MIAs: LiRA [5], Attack R [45] and RMIA [51]. Impressively, when measured on predicting vulnerability to at least one of these attacks, our method performs on par with Attack R and RMIA with a few shadow models, despite incurring a fraction of the computational cost.
We then explore other loss trace aggregation methods - all of which analyze the per-sample training dynamics captured by the per-epoch loss trace. We find that multiple aggregation methods show strong performance on par with LT-IQR, with slight variability depending on the dataset. Additionally, we perform an ablation study on LT-IQR’s quantile parameters, showing the method to be quite robust to specific parameter values.
Finally, we analyze the distribution of LT-IQR vulnerability scores throughout training, and note the distinctive changes in the distribution shape as the training progresses and the model memorizes more. This suggests that LT-IQR could also potentially be used to estimate model-level vulnerability in addition to detecting vulnerable points, as considered before.
@@
Our cheap yet effective approach enables model developers to assess the privacy risks associated with a given model release and deploy appropriate mitigations if needed, including unlearning, data removal, or retraining with differential privacy. Given its low computational cost, our method could be especially useful to compare the privacy leakage of different hyperparameter setups, or as an early indication to run a full privacy attack on a subset of candidate models.
@@ 2.1 Membership Inference Attacks
Membership Inference Attacks (MIAs) aim to determine whether a certain data sample was used to train a given model. They typically rely on a strong yet realistic attacker [36, 33], and are used by model developers as an upper bound for other attacks including extraction  [8, 35, 28] and attribute inference. In many cases the vulnerability to an MIA does not pose a high risk in practice: either because a realistic attacker would not have access to the full user record, or because membership alone does not imply the possession of a sensitive attribute. However, as MIA is the easiest type of attack for an adversary to perform, it is often used to estimate empirical privacy risk, under the assumption that protection against an MIA implies the protection against all other types of attack [36].
MIAs are typically evaluated with TPR at low FPR - a metric measuring an attacker’s ability to confidently and correctly predict membership of a small number of points, irrespective of the attack’s performance on average. This recognizes the fact that in a privacy setup, an attacker making confident predictions about a small minority of points possesses higher risk than an attacker making predictions slightly better than random on every point.
@@
Existing low-cost attacks [46, 51] either compromise utility relative to shadow-model-based approaches or still necessitate training some shadow models [51], which remains computationally expensive given the positive correlation between attack performance and the number of shadow models employed.
@@ 2.2 Individual vulnerability
An essential step in DP-SGD [1] (a differentially private variation of SGD) is clipping the per sample gradient to bound the sensitivity of the output to any individual sample. This demonstrates the disparate nature of individual data points’ contribution to the model - where clipping is necessary to ensure the model doesn’t learn too much from any one sample.
@@
It then naturally follows, that in the absence of gradient clipping, some points would be memorized more than others. Empirical privacy risk is indeed typically concentrated around a small number of data outliers [13, 3, 7].
Specifically from the perspective of MIA vulnerability, two prior works explore the correlation between the attack score and certain data properties. Leemann et al. [21] investigate the connection between MIA score and certain per-instance properties: final loss, parameter gradients, local loss Hessian, and others, identifying the final loss to be the best predictor of the MIA performance. They conclude that the loss is the best predictor of the MIA vulnerability. Tobaben et al. [43] study the impact of the average gradient norm throughout training on the MIA performance and find a strong correlation across multiple datasets. Unlike our work, however, both Leemann et al. [21] and Tobaben et al. [43] study the correlation between the data properties and attack success while we focus on the few most vulnerable samples and using training artifacts as a source of information.
@@ 2.3 Tracing memorization
Unlike privacy analyses described above, we consider an artifact-based approach where the model developer has access to all intermediate model states during training. This setup allows us to use the loss throughout training, for instance, as an additional source of information on top of the commonly used final loss.
@@
Some prior works have considered tracing privacy metrics throughout training. Pruthi et al. [37] tracks the loss before and after each training step to estimate the influence of a training sample on model’s predictions. Lesci et al. [23] characterize a model’s counterfactual memorization [52] by observing losses on a subset of data throughout training. Liu et al. [26] and Li et al. [24] propose an MIA where an attacker performs model distillation and tracks the loss throughout the distillation process.
@@
For differentially private training, privacy auditing often adopts a similar threat model to us. Privacy auditing has emerged as a way to either assess the tightness of the theoretical privacy analysis [16, 35, 50] or verify the correctness of DP algorithm implementations [44, 34]. Privacy auditing empirically estimates the lower bound for the differential privacy guarantees of the model training process by running privacy attacks and measuring their success. Assumed threat models can range from a weak adversary with only black-box access to the model [16, 35, 41], to a white-box adversary capable of observing intermediate model updates [35, 34, 41], to a very strong adversary with full control over the training dataset apart from the target record [35].
@@
Proposed attack methods vary depending on the threat model. For black-box attacks, a simple loss threshold [46] is typically used. For white-box adversaries with access to the intermediate model states, a threat model similar to the one we assume in this work, prior works suggested using aggregate metrics over the entire training process. Proposed methods include taking the maximum or average of the loss [35], or computing the dot product between the canary gradient and the overall batch update [34, 41].
As the goal of the privacy auditing is to provide a tight estimate of the differential privacy guarantees, prior works typically rely on data poisoning attacks, instantiating the MIAs against data samples (canaries) purposely injected into the training dataset. This allows the auditor to construct the worst possible scenario with the maximum privacy leakage. In contrast, we here focus on estimating the privacy risks of realistic data points in a non-adversarially constructed dataset.
3 Task
In this paper we set the task of identifying the most vulnerable points among the training samples. Given the training dataset, we aim to rank the data points by their vulnerability, providing a ranked list as an output. As the privacy risk is typically concentrated in a small proportion of the training dataset - and only this subset is highly relevant for any potential mitigation - we focus our evaluation on the top end of the output ranked list.
For evaluation purposes we assume access to binary ground truth vulnerability labels representing actual attack results (which we expect to be highly imbalanced). Given these labels, we evaluate our vulnerability detection methods using Precision@k and Recall@k, which measure what fraction of the k points with highest vulnerability scores are actually vulnerable. For a vulnerability scoring function $\phi$ and a set of vulnerable points $V$, we define Precision@k as:
[…]
where $\text{top}_{k}(\phi)$ represents the k points with the highest vulnerability scores according to $\phi$. In other words, if we sort all training points by their vulnerability score $\phi(x,y)$ in descending order and take the first k points, Precision@k tells us what fraction of these points are actually vulnerable.
Similarly, we define Recall@k as:
[…]
Here, recall measures the fraction of all vulnerable points we successfully identified within our top k predictions.
For acquiring the ground truth labels, we perform one or multiple state-of-the-art shadow model-based MIAs, characterized by their ROC curves. We consider a fixed and low FPR threshold $\alpha$ (typically $\alpha=10^{-3}$) and define samples that are identified as members at this strict threshold as vulnerable. We use the terms “point” and “sample” interchangeably. For a given $\alpha$, we select the threshold $\tau_{\alpha}$ such that $\text{FPR}(\tau_{\alpha})=\alpha$.
Definition 3.1 (Vulnerable points).
We initially consider vulnerability to LiRA, as it is the best performing attack with the highest number of vulnerable points - a clear choice for a hypothetical attacker. Additionally, we show that most points vulnerable to other attacks are also vulnerable to LiRA, but not the other way around. Only about 7% of points vulnerable to RMIA and Attack R are not covered by LiRA (see Fig. 4).
We also note that perfect Recall@k is often not achievable due to the relationship between k and the total number of vulnerable points identified at a given FPR threshold. For instance, with $k=1\%$ (250 samples) and 2435 total vulnerable points in our standard setup, the maximum theoretical recall is only 10%. We report the maximum achievable Recall@k values for each experimental configuration and discuss this property further in Appendix B.
@@ 4.1.1 Model training
Let $f_{\theta}:\mathcal{X}\rightarrow[0,1]^{N}$ be a machine learning classification model with an input domain $\mathcal{X}$ and an output domain $\mathcal{Y}$ with $N=|\mathcal{Y}|$ classes, parametrized by a parameter vector $\theta$. $f_{\theta}(x)$ outputs a probability distribution over $N$ classes. For brevity, we denote the probability of a given class $y$ as $f_{\theta}(x)_{y}$. We denote train and test datasets as $D_{\text{train}}=\{(x_{i},y_{i})\mid x_{i}\in\mathcal{X},y_{i}\in\mathcal{Y}\}$ and $D_{\text{test}}=\{(x^{*}_{i},y^{*}_{i})\mid x^{*}_{i}\in\mathcal{X},y^{*}_{i}\in\mathcal{Y}\}$, respectively. Let $\mathcal{T}$ be a training process such that the final model weights $\theta$ are obtained as $\theta\leftarrow\mathcal{T}(D_{\text{train}})$. We consider the training process to be performed over $S$ epochs, and additionally denote the intermediate model state after $s=1\dots S$ epochs as $\theta_{s}\leftarrow\mathcal{T}_{s}(D_{\text{train}})$. Note that $\theta_{0}$ is the initial model state and $\theta_{S}=\theta$ is the parameter vector of a fully trained model.
@@
Specifically for our method, we denote the loss on a sample $(x,y)$ after $s$ training epochs as $\ell_{s}(x,y)=\ell(f_{\theta_{s}}(x),y)$. We also define $\mathcal{L}_{\text{train}}^{s}$ and $\mathcal{L}_{\text{test}}^{s}$ as average losses at epoch $s$ on train and test sets respectively.
@@
4.2 Membership Inference Attacks
A membership inference attack (MIA) aims to determine whether a given data sample was used to train a model. Formally, we define a membership inference attack as a function $A:\mathcal{X}\times\mathcal{Y}\times\Theta\rightarrow[0,1]$, where $A(x,y,\theta)=1$ indicates that the attacker believes the sample $(x,y)$ was used to train the model with parameters $\theta$, and $A(x,y,\theta)=0$ indicates when they believe it was not.
In particular, we consider SOTA MIAs that output a confidence score, where higher values indicate stronger belief that the sample was used during training. The binary classification decision is then made by comparing this score against a threshold $\tau$: